OpenStudy (anonymous):

j^3 = 61 + i^6 Find all the integer solution of i and j.

OpenStudy (anonymous):

One is i=2 and j=5

OpenStudy (anonymous):

What about other?

OpenStudy (anonymous):

\[j^3-(i^2)^3=61\] \[(j-i^2)(j^2+ij+i^2)=61\] 61 is prime so 61(1) are the only factors hence \[j-i^2=1,j^2+ij+i^2=61\]

OpenStudy (anonymous):

Looks like 4 solution.

OpenStudy (anonymous):

but we imgore the fact that it can be vise versa 1(61)

OpenStudy (anonymous):